You are provided the following information related to the Three Stooges Investment Funds LTD.

Portfolie Metrics Larry Curly Moe
Total Return 10% 14% 8%
Beta 0.9 1.5 0.8
Standard Deviation 12% 30% 10%
Correlation 0.3 0.6 0.4
Risk-Free Rate 4% 4% 4%
Equity Risk Premium 8.5% 8.5%

8.5%

AAA Bond Rate 9.5% 9.5% 9.5%

Based on the above data, calculate the Sharpe Ratio for the Larry and Curly funds. How do they compare (explain what the ratio is measures)? Please show all work and use at least 4 decimal places.
Answer #1

Sharpe ratio = (total return - risk free rate)/ standard deviation

ratio measure return on per unit risk. Higher the vlaue better for the stock.

Sharpe ratio for Larry = (10 - 4)/12 = .5

Sharpe ratio for Curly = (14 - 4) /30 = .3333

since the vlaue of sharpe ratio for Larry is greater it is generating more return per unit of risk and its performance is better than Curly per unit of risk.

Standard deviation measure the risk of the stock.

Please rate the answer.